# An End-to-End Room Geometry Constrained Depth Estimation Framework for Indoor Panorama Images

### v1

@@ III-B Room Geometry-based Background Depth Resolving Strategy
After determining the camera height and the distance between the camera and the ceiling, we can use the wall information from the room layout predictions to calculate the corresponding background depth maps for each pixel belonging to the background wall. The depth maps for the ceiling and floor can be simply calculated based on the angles calculated from the pixels using the following formula:
$d^{i}_{floor}=\frac{|PB|\times H}{(u^{i}_{floor}-0.5H)\times\pi}$ (4)
$d^{i}_{ceil}=\frac{|AP|\times H}{(0.5H-u^{i}_{ceil})\times\pi}$
@@
III-C Background-segmentation-guided Fusion Mechanism
@@
As mentioned above, based on the corase depth map and the background depth map, our framework needs to fuse the predicted depth $S^{p}_{depth}$ and the background depth $S_{back}$ to obtain an accurate final depth estimation result $S^{final}_{depth}$. Considering that $S_{back}$ is naturally suitable as an upper bound for depth estimation when the layout prediction is relatively accurate. Then $S_{back}$ as an upper bound can be used to ensure that the depth of pixels originally belonging to the rescue, ceiling, ground and other areas do not exceed the area of the room itself. The depth value of the pixel belonging to the foreground object area can be as close to $S^{p}_{depth}$ as possible. Therefore, in the process of fusing the two depth maps, we need a weight to determine whether the current pixel is an object belonging to the foreground area or the background area. The probability map $S_{seg}$ predicted by the background-segmentation task in our framework can just be used as such a fusion weight. Therefore, the entire fusion process can be described as the following formula:
$S^{final}_{depth}$ $=\{d^{final}_{ij}|i=1,2,.,W;j=1,2,.,H\};$ (6)
$etl.$ $\ d^{final}_{ij}=d^{back}_{ij}\times p_{ij}+d^{p}_{ij}\times(1-p_{ij});$
$p_{ij}\in S_{seg};$
$d^{back}_{ij}\in S_{back};$
$d^{p}_{ij}\in S^{p}_{depth};$
III-D Objective Function
[…]
When training the layout branch, we refer to the settings of HorizonNet and use the binary cross entropy and L1 loss function to calculate. When training the depth estimation branch, we used the same Huber[43] (or Berhu[44]) loss and gradient loss[45] as the baseline Panoformer. The hyper-parameters within the objective functions of these two branches were also consistent with those of HorizonNet and PanoFormer.
Considering that existing panoramic image datasets usually do not contain background segmentation annotations. When training the background segmentation decoder, we first use the background depth map $S_{back}$ predicted by the framework and the corresponding ground-truth depth map $S^{gt}_{depth}$ to calculate the corresponding background segmentation label $S^{gt}_{seg}$. The specific solution formula is as follows:
$D^{res}=|S^{gt}_{depth}-S_{back}|;$ (7)
$S^{gt}_{seg}=\{p^{gt}_{ij}=\Gamma(d^{res}_{ij}<\gamma)|d^{res}_{ij}\in D^{res}\}$
where $\Gamma$ is a binary function, which takes the value 1 when the condition in the brackets is true and takes the value 0 when it is false.
When optimizing the partial network for the background-segmentation task, we use a simple focal loss as our object function, which is calculated as follows:
$L_{seg}=-\frac{1}{M}\sum^{M}_{u=1}\frac{1}{H*W}\sum^{W}_{i=0}\sum^{H}_{j=0}\alpha(1-\hat{p}_{ij})^{\eta}log(\hat{p}_{ij});$ (8)
where $\hat{p}_{ij}=p_{ij}$ if $p^{gt}_{ij}=1$ else $\hat{p}_{ij}=1-p_{ij}$. Following the common setting of focal loss, we set the $\alpha$, $\beta$ as 0.5, 2.0 relatively.
In summary, the overall object function of our proposed framework can be described as:
$L_{all}=\lambda_{1}*L_{layout}+\lambda_{2}*L_{depth}+\lambda_{3}*L_{seg}$ (9)
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are three manual setting parameters.
@@ IV-A Experiment Setting
In this section, we introduce the experiments to verify the proposed depth estimation framework. In our experiments, we selected three datasets: Stanford2D-3D[46], MatterPort3D[47], and Structure3D[48]. Stanford2D-3D and MatterPort3D are two datasets collected in the real world, while the Structure3D dataset is a synthesized dataset. Real-world datasets contain a large amount of noise points introduced during the acquisition process as described in OnimiDepth[13]. This noise introduces local jumps that severely hinder model performance. Therefore, before training models on Stanford2D3D and Matterport3D, we used a proposed layout-based dataset denoise strategy to constrain the depth of these noise points to within the room.
Stanford2D-3D: The Stanford2D-3D dataset contains 1413 panoramic images collected from 3 types of buildings divided into 6 large areas. We follow the official practice to divide the dataset into training and test sets, and downsample all depth maps and RGB images to $512\times 1024$ size images.
@@
Metric: Following the previous works, we use some standard evaluation metrics, which include: relative error (abs rel), squared relative error (sq rel), root mean squared error (rmse), and three threshold percentage $\delta<\varsigma^{t}(\varsigma=1.25,t=1,2,3)$ denoted as $\delta_{t}$.
Dataset denoise strategy: Considering that datasets collected in the real world contain a large amount of noisy points, we have also designed a denoising strategy for panoramic depth map datasets. As mentioned earlier, our method uses the outputs of the layout decoder and depth decoder to calculate the corresponding background depth map. Here, we use this calculated background depth map to constrain the dataset collected in the real world. First, for areas in the real-world depth ground-truth maps where measurement fails, we directly replace them with the background depth value. Then, our method sets a threshold to judge the ground-truth depth. If the depth point in the ground-truth dataset of the original data is converted to 3D space and its coordinates are outside the room and the distance from the room wall is more than 1 meter, we identify the pixel as a noise pixel. For these pixels, we also use the depth in the background depth map as a replacement.
@@
Parameter setting: The training objective function used by our framework is shown in Section 4.4, which contains three subcomponents: $L_{front}$, $L_{depth}$, and $L_{layout}$. Considering that some datasets do not contain complete layout annotations, and our framework is extremely dependent on the accuracy of layout estimation, we will first extract all layout annotations from the three selected datasets to form a large layout dataset to pretrain the layout task. Then, when training the depth estimation task on specific datasets, we decide whether to finetune the weights of the layout and feature encoder parts based on whether the dataset contains complete layout annotations. Specifically, the Stanford2D-3D data does not provide complete layout annotations, so we set the corresponding hyper-parameter $\lambda_{1}$ to 0 when training our framework, and the corresponding $\lambda_{2},\lambda_{3}$ to 1.0 and 0.4. The Matterport3D and Structure3D datasets provide complete layout annotations, so we set $\lambda_{1},\lambda_{2},\lambda_{3}$ to 0.01, 1.0, and 0.4, respectively. In addition, some hand-tuned hyper-parameter settings related to this study, such as $\rho_{1}$ and $\rho_{2}$ in equation was set to 1.0 and 0.5. Other hyper-parameters that are not relevant to this study are set according to the common specifications in the current field.
IV-B Performance Comparison
Figure 5: The 3D visualize results of final depth estimation. For each scene, we selected three perspectives: top view, side view, and internal perspective to display the three-dimensional visualization effect of the point cloud converted from the depth map.
IV-B1 Quantification Comparison with SoTA Depth Estimation Methods
@@
Matterport3D EGFormer[11] ICCV 2023 0.8158 0.9390 0.9735 0.6025 0.1517 0.1473
OminiFusion[19] CVPR 2022 0.9040 0.9757 0.9919 0.4261 0.0552 0.0900
Bifuse[12] CVPR 2020 0.8452 0.9319 0.9632 0.6295 0.2408 0.3470
UniFuse[28] IEEE RAL 2021 0.8897 0.9623 0.9831 0.4941 - 0.2814
@@
RSDNet[50] CVPR 2022 0.443 0.097 0.248 0.906 0.971 -
@@
Ours 0.9199 0.9820 0.9984 0.2436 0.0425 0.0829
TABLE II: Quantification comparison with State of The Art depth estimation methods on Stanfor2d3d
@@
Stanford2d3d EGFormer[11] ICCV 2023 0.8185 0.9338 0.9736 0.4974 0.1408 0.1528
OminiFusion[19] CVPR 2022 0.8940 0.9714 0.9900 0.3715 0.0543 0.0961
Bifuse[12] CVPR 2020 0.8660 0.9580 0.9860 0.4142 0.1209 0.2343
UniFuse[28] IEEE RAL 2021 0.8711 0.9664 0.9882 0.3691 - 0.2082
HRDFuse[29] CVPR 2023 0.8941 0.9778 0.9923 0.3452 0.0503 0.0984
FreDSNet[52] 2022 0.8424 0.9583 0.9863 0.2727 0.0952 0.1327
RSDNet[50] CVPR 2022 0.394 0.098 0.209 0.903 0.974 -
SN360[51] IEEE Access 2025 0.2917 - - 0.9369 0.9846 0.9942
GLPanoDepth[23] IEEE TIP 2024 0.9015 0.9793 0.9901 0.3493 - 0.1932
Panoformer[10] ECCV 2022 0.9394 0.9838 0.9941 0.3083 0.0405 0.0619
SGFormer[15] IEEE TCSVT 2025 0.8998 0.9693 0.9908 0.3406 - 0.2017
Ours 0.9479 0.9857 0.9943 0.2359 0.0285 0.058
TABLE III: Quantification comparison with State of the Art depth estimation methods on Structured3D
@@
Structure3d EGFormer[11] ICCV 2023 0.7979 0.9071 0.9455 0.6841 0.4509 0.2205
OminiFusion[19] CVPR 2022 0.6921 0.8831 0.9501 0.4951 - 0.2981
Bifuse[12] CVPR 2020 0.8594 0.9400 0.9672 0.5213 0.2455 0.1573
UniFuse[28] IEEE RAL 2021 0.8542 0.9399 0.9676 0.5016 0.2319 0.1506
HRDFuse[29] CVPR 2023 0.7561 0.9161 0.9631 0.4061 - 0.2451
BGDNet[32] CVPR 2024 0.8336 0.9377 0.9731 0.3490 - 0.1656
SGFormer[15] IEEE TCSVT 2025 0.9613 0.9896 0.9957 0.2429 - -
PanoFormer[10] ECCV 2022 0.8943 0.9536 0.97431 0.3017 0.1201 0.1546
Ours 0.9679 0.9907 0.9983 0.1935 0.0414 0.0613
Figure 6: Visualization of background depth map guided dataset denoising effect diagram.
[…]
The Stanford2d3d dataset is a relatively early dataset, and various existing depth estimation methods have achieved relatively good performance on this dataset. However, on this dataset, the noise processing strategy and multi-task depth estimation framework proposed in this paper still show excellent effectiveness. Specifically, our method significantly outperforms the current SOTA method by 3.68 percentage points in the RMSE, the evaluation indicator that the industry attaches the most importance to.
At the same time, in MRE and MAE, which are both numerical indicators, the performance advantage of our method over the existing methods is not as obvious as RMSE, but it is still intuitive, which further proves the robustness and superiority of the method proposed in this paper. In the three indicators $\delta_{1},\delta_{2}$ and $\delta_{3}$ at the ratio level, our method also has a leading advantage, but this advantage is not as obvious as the numerical indicators. We believe that this is because the existing methods have reached a very high level in the three indicators at the ratio level, so it is more difficult to achieve higher indicators.
@@
On the Matterport3D dataset, our method also achieved the current best performance in RMSE, the core evaluation indicator of depth estimation, which fully demonstrated the effectiveness of the proposed method and data processing strategy in key accuracy metrics. However, on the relative error threshold indicator ($\delta<$ 1.25), our method failed to achieve the best results. We believe this is related to the introduced layout constraints: this constraint tends to optimize the absolute error between the prediction and the true value, but in areas with small true depth values, this optimization may result in a limited range of change in the predicted value, making it more difficult to meet the strict 1.25 times relative error requirement. Despite this, we still achieved a level close to SOTA on this indicator. Importantly, this method is significantly ahead of existing work in both the core rmse and rmslog indicators.

### v2

@@ III-B1 Camera Height Optimize
Through the above process, we can obtain a set of camera heights and distances from the camera to the ceiling. However, this calculation process is greatly affected by depth estimation errors. To reduce this impact, we use the horizontal line passing through point $O$ as a reference and perform the above calculation process at each occlusion-free point $P$ on this line segment. Finally, we average all the calculated camera heights to obtain the final $\tilde{h}_{c}$. Ultimately, averaging $\hat{h}_{c}$ and $\tilde{h}_{c}$ yields a relatively accurate camera height $h_{c}$.
III-B2 Enclosed Background Depth Resolving
After determining the camera height $h_{c}$ and the distance between the camera and the ceiling, we can use the wall information from the room layout predictions to calculate the corresponding background depth maps for each pixel belonging to the background wall. The depth maps for the ceiling and floor can be simply calculated based on the angles calculated from the pixels using the following formula:
$d^{i}_{floor}=\frac{h_{c}\times H}{(u^{i}_{floor}-0.5H)\times\pi},$ (3)
$d^{i}_{ceil}=\frac{|AP|\times H}{(0.5H-u^{i}_{ceil})\times\pi},$
@@ III-B2 Enclosed Background Depth Resolving
III-C Fusion Mask Generation Component
A crucial step in constraining the depth decoder’s prediction using the calculated background depth is identifying which regions in the input panoramic image are valid. Previous work, BGDNet, simply used the semantic segmentation mask output by the SAM model as the criterion for this. This approach effectively distinguishes which parts of a room’s interior are foreground objects and which areas belong to the background walls. However, this approach relies on two prerequisites: 1) no large areas in the input image extend beyond the room structure; and 2) the room structure itself is regular and conforms to the Manhattan layout. Previous work has mentioned that any irrationally shaped room structure can essentially be divided into regular and irregular regions. The room layouts typically provided by datasets describe enclosed regular regions within the room structure. Therefore, to make the background depth-based method more generalizable, we need to determine not only whether a pixel belongs to the foreground object but also whether the region belongs to the regularly enclosed region defined by the room layout.
In our framework, we treat the distinction between foreground and background regions, and between regular enclosed regions and irregular regions, as two independent binary semantic segmentation tasks. Both of these semantic segmentation tasks are performed by our region segmentation decoder. This decoder outputting the irregular region mask $S_{ir}$ and the background mask $S_{seg}$. In $S_{ir}$, each element $s^{ir}_{uv}$ represents the probability that the pixel at row $u$ and column $v$ in the input panorama belongs to an irregular region. Similarly, in $S^{seg}$, each element $s^{seg}_{uv}$ denotes the probability that the pixel belongs to the background. To properly fuse coarse depth prediction and background depth, we need to calculate a fusion weight to balance their respective proportions in the final output. Here, we first binarize each element in the region mask $S_{ir}$ using a threshold of $0.5$. Then, we use the sigmoid function to compress the value of each element in $S_{seg}$ to the range $[0,1]$. Finally, we multiply each element in $S_{ir}$ by each element in $S_{seg}$ to obtain the final fusion weight mask $S_{weight}$.
III-D Adaptive Fusion Component
@@
As mentioned above, based on the depth decoder’s prediction and the enclosed background depth map, our framework needs to fuse the predicted depth $S^{p}_{depth}$ and the enclosed background depth $S_{back}$ to obtain an accurate final depth estimation result $S^{final}_{depth}$. Considering that $S_{back}$ is naturally suitable as an upper bound for depth estimation when the layout prediction is relatively accurate.Then $S_{back}$ as an upper bound can be used to ensure that the depth of pixels originally belonging to the rescue, ceiling, ground and other areas belonging to regular regions do not exceed the area of the room itself. And, the depth value of the pixel belonging to the foreground object or irregular region can be as close to $S^{p}_{depth}$ as possible. The entire fusion can be described as following formula:
$S^{final}_{depth}$ $=\{d^{f}_{uv}|u=1,2,.,W;v=1,2,.,H\};$ (5)
$etl.$ $\ d^{f}_{uv}=s^{uv}_{weight}\times d^{b}_{uv}+(1-s^{uv}_{weight})\times d^{p}_{uv},$
$s^{uv}_{weight}\in S_{weight};d^{b}_{uv}\in S_{back};d^{p}_{uv}\in S^{p}_{depth},$
III-E Objective Function
Our framework as a whole includes four task decoders: layout estimation, depth estimation, camera pose estimation, region segmentation. The objective functions corresponding to the four decoders can be expressed as $L_{layout}$, $L^{p}_{depth}$, $L_{region}$, and $L_{pose}$, respectively. Since most existing datasets lack comprehensive annotations for room layout estimation, we first aggregate all layout-related annotations from the Structured3D and Matterport3D datasets to form a large-scale layout dataset. Following the practice described in[31], we apply a simple script to convert the existing room layout annotations—which may contain occlusions or ambiguities—into regular and enclosed layouts. Using this generated dataset, we first pre-train the room layout branch of our proposed framework. For this stage, similar to HorizonNet, we adopt a combination of binary cross-entropy loss and L1 loss. For the depth estimation branch, we follow Panoformer and employ a Huber[9] loss together with a gradient loss[26], keeping the hyper-parameter settings consistent with those in the original HorizonNet[28] and Panoformer papers respectively. For the two semantic segmentation subtasks, we use the commonly adopted binary cross-entropy loss and Dice loss[17]. As for the camera pose estimation, we simply apply a smooth L1 loss. Overall, the total loss function of our proposed framework can be expressed as:
$L_{all}=\lambda_{1}*L_{layout}+\lambda_{2}*L_{depth}+\lambda_{3}*L_{region}+\lambda_{4}*L_{pose},$ (6)
where $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ are four manual setting parameters.
@@ IV-A Experiment Setting
In this section, we introduce the experiments to verify the proposed depth estimation framework. In our experiments, we selected three datasets: MatterPort3D[5], Structure3D[43], Replica[30]. MatterPort3D and Reiplica is collected in the real world, while the Structure3D dataset is a synthesized dataset.
@@
Replica: The Replica dataset is captured using an RGB-D camera rig with an IR projector from real-world indoor scenes. We follow the BGDNet[6] render 3554 images from these 3D models by setting camera in random position on the floor.
Metric: Following the previous works[3, 25, 40], we use some standard evaluation metrics to measure our methods, which include: mean relative error (MRE), mean average error (MAE), root mean squared error (RMSE), and three threshold percentage $\delta<\varsigma^{t}(\varsigma=1.25,t=1,2,3)$ denoted as $\delta^{t}$. For the room layout task, we used 2D IoU, 3D IoU, RMSE, and $\delta^{1}$ for evaluation. For the camera pose task, we used relative rotation error(RRA), relative translation error(RTA), and area under the curve(AUC) for evaluation, with a threshold of 30 for each metric. For the region segmentation task, we used mIoU for evaluation.
@@
Parameter setting: The training objective function used by our framework is shown in Section III. B , which contains four subcomponents: $L_{pose}$, $L_{region}$, $L_{depth}$, and $L_{layout}$. During pre-training of the layout decoder, we set the $\lambda_{1}$ to 0 and other parts to 0, while during the training the training of the other decoders, we set $\lambda_{2},\lambda_{3},\lambda_{4}$ as 1.0, 0.4, 0.6. Other hyper-parameters that are not relevant to this study are set according to the common specifications in the current field.
@@
Matterport3D Panoformer[25] ECCV 2022 0.9184 0.9804 0.9916 0.3635 0.0571 0.1013
EGFormer[38] ICCV 2023 0.8158 0.9390 0.9735 0.6025 0.1517 0.1473
@@
GLPanoDepth[3] IEEE TIP 2024 0.8641 0.9561 0.9808 0.5223 - 0.2998
GADFNet[13] IEEE TCSVT 2025 0.9047 0.9668 0.9868 0.4564 0.0963 0.2592
@@
DepthAnyPanorama[20] arxiv 2025 0.8518 0.9113 0.9764 0.7510 0.4456 0.7845
DepthAnyDirection[16] ICLR 2026 0.9561 0.9860 0.9982 0.2882 0.0766 0.0912
Ours 0.9399 0.9852 0.9984 0.2236 0.0411 0.0629
TABLE II: Quantification comparison with State of the Art depth estimation methods on Structured3D
@@
Structure3d PanoFormer[25] ECCV 2022 0.8943 0.9536 0.97431 0.3017 0.1201 0.1546
EGFormer[38] ICCV 2023 0.7979 0.9071 0.9455 0.6841 0.4509 0.2205
HRDFuse[1] CVPR 2023 0.7561 0.9161 0.9631 0.4061 - 0.2451
SGFormer[40] IEEE TCSVT 2025 0.9613 0.9896 0.9957 0.2429 - -
GADFNet[13] IEEE TCSVT 2025 0.8647 0.9468 0.9899 0.3524 0.0812 0.1944
DepthAnyPanorama[20] arxiv 2025 0.8918 0.9243 0.9813 0.6641 0.4456 0.5125
DepthAnyDirection[16] ICLR 2026 0.9361 0.9713 0.9912 0.2914 0.0686 0.0992
BGDNet[6] CVPR 2024 0.8336 0.9377 0.9731 0.3490 - 0.1656
Ours 0.9679 0.9907 0.9983 0.1935 0.0414 0.0613
TABLE III: Quantification comparison with State of the Art depth estimation methods on Replica
@@
Replica HRDFuse[1] CVPR 2023 0.7711 0.9221 0.9541 0.3891 - 0.1421
DepthAnyPanorama[20] arxiv 2025 0.9012 0.9335 0.9898 0.4432 0.2121 0.3132
DepthAnyDirection[16] ICLR 2026 0.9431 0.9668 0.9933 0.2712 0.0733 0.1011
BGDNet[6] CVPR 2024 0.8554 0.9365 0.9624 0.3456 - 0.1678
Ours 0.9233 0.9866 0.9992 0.2101 0.0512 0.0629
IV-B Region Segmentation Decoder’s Annotation Generation
IV-B1 The generation process of background segmentation
To efficiently fuse the predictions from the background and depth decoders, we first need information to guide the model in identifying which regions belong to the background and which to the foreground. For region segmentation, we refer to BGDNet’s approach, using the corresponding semantic segmentation results to uniformly set all ground, ceiling, and wall surfaces in the scene as the background region. Unlike BGDNet, which uses SAM to segment the input image, we directly use the semantic segmentation annotations provided by the dataset to generate the corresponding masks.
IV-B2 The generation process of irregular region segmentation
As mentioned in the main text, the background depth generated by BGDNet often ignores the fact that in some scenes, the depth threshold may exceed the background space defined by the room layout, or the entire indoor scene itself may not be a regular Manhattan layout. However, non-Manhattan layouts or irregularly shaped rooms will always intrude a regular enclosed Manhattan region, and the room layout annotations provided in the dataset are usually this regular enclosed Manhattan region. Therefore, to make this approach of fusing background and depth decoders more generalizable, we need information to help the model distinguish which regions belong to the regular enclosed region.
Here, to generate a reasonable irregular region segmentation mask, we first calculate the scene background depth for each panel input image using the room layout annotations and camera height from the entire dataset. Then, we use the generated background segmentation mask annotation to select the depth of the fused region from both the background depth and ground truth depth annotations. Finally, we calculate the difference between the depths of the fused regions selected from the two depths and apply a threshold to obtain the corresponding irregular region segmentation mask annotation. In the process of creating the binary mask, we use a threshold of 0.05, which is based on our specific observations in the experiments.
IV-C Performance Comparison
IV-C1 Quantification Comparison with SoTA Depth Estimation Methods
Figure 4: The 3D visualize results of final depth estimation. For each scene, we selected three perspectives: top view, side view, front view, and internal perspective to display the three-dimensional visualization effect of the point cloud converted from the depth map.
In this section, we will conduct performance comparison experiments on three datasets: Matterport3d, Structured3d, and Replica. The results have been shown in Table I- III . To validate the effectiveness of our approach, we compare our RGCNet with the current state-of-the-art methods[38, 18, 32, 14, 1, 3, 25, 40, 6, 4, 11, 21], including strategies for bi-projection fusion, long-range dependencies, multi-task learning, and background-based methods.
@@ IV-C1 Quantification Comparison with SoTA Depth Estimation Methods
On the Matterport3D dataset, our method also achieved the current best performance in RMSE, the core evaluation indicator of depth estimation, which fully demonstrated the effectiveness of the proposed method and data processing strategy in key accuracy metrics. However, on the relative error threshold indicator ($\delta<$ 1.25), our method failed to achieve the best results. We believe this is related to the introduced layout constraints: this constraint tends to optimize the absolute error between the prediction and the true value, but in areas with small true depth values, this optimization may result in a limited range of change in the predicted value, making it more difficult to meet the strict 1.25 times relative error requirement. Despite this, we still achieved a level close to SOTA on this indicator. Importantly, this method is significantly ahead of existing work in RMSE.